# HI-SLAM2: Geometry-Aware Gaussian SLAM for Fast Monocular Scene Reconstruction

### v2

@@ Abstract
We present HI-SLAM2, a geometry-aware Gaussian SLAM system that achieves fast and accurate monocular scene reconstruction using only RGB input. Existing Neural SLAM or 3DGS-based SLAM methods often trade off between rendering quality and geometry accuracy, our research demonstrates that both can be achieved simultaneously with RGB input alone. The key idea of our approach is to enhance the ability for geometry estimation by combining easy-to-obtain monocular priors with learning-based dense SLAM, and then using 3D Gaussian splatting as our core map representation to efficiently model the scene. Upon loop closure, our method ensures on-the-fly global consistency through efficient pose graph bundle adjustment and instant map updates by explicitly deforming the 3D Gaussian units based on anchored keyframe updates. Furthermore, we introduce a grid-based scale alignment strategy to maintain improved scale consistency in prior depths for finer depth details. Through extensive experiments on Replica, ScanNet, Waymo Open, ETH3D SLAM and ScanNet++ datasets, we demonstrate significant improvements over existing Neural SLAM methods and even surpass RGB-D-based methods in both reconstruction and rendering quality. The project page and source code will be made available at https://hi-slam2.github.io/.
Index Terms:
Visual SLAM, Dense Reconstruction, Deep Learning for Visual Perception.
@@ III-A Comparison to HI-SLAM
  • Hierarchical Optimization: Our system employs a multi-stage optimization pipeline that includes online tracking with local BA, online loop closure with $Sim(3)$-based PGBA, and global full BA. Last but not least, the joint refinement of both Gaussian map parameters and camera poses, which couples the tracking frontend with the mapping backend more tightly. In contrast, HI-SLAM only performs pose optimization in the tracking frontend, which can lead to inconsistencies between estimated poses and map geometry. This hierarchical approach reduces the Absolute Trajectory Error (ATE) by 29.3% compared to online tracking alone (Table IX) on the Replica dataset, yielding a globally more consistent reconstruction.
@@ IV-D Camera Tracking Accuracy
We evaluate the camera tracking accuracy of our system against state-of-the-art dense visual SLAM methods on indoor Replica, ScanNet, ETH3D SLAM datasets and outdoor Waymo datasets, including comparisons with RGB-D based approaches. The ATE results in Tables I and II demonstrate the superior tracking accuracy of our system. RGB-D methods SplatTAM [31] and MonoGS [32], despite having access to depth measurements for map-based tracking, achieve lower accuracy than hybrid approaches. Our system, along with Splat-SLAM [33], represents the class of hybrid methods that effectively combine dense SLAM with deep learning foundations. The global BA of DROID-SLAM [20] was enabled in all experiments. While it employs global BA, our additional global pose and map joint refinement further improves tracking accuracy beyond the baseline method. We further compare our system on the ETH3D SLAM dataset with state-of-the-art sparse and dense methods, including ORB-SLAM3 [17], DPVO [73], DPV-SLAM [74], and MASt3R-SLAM [75]. Figure 6 illustrates the cumulative success curves based on ATE thresholds. As none of the methods can successfully track all sequences, we use the Area Under the Curve (AUC) metric to assess both accuracy and robustness with an upper ATE threshold of 0.5m. Our system achieves the highest AUC among all methods. Out of total 61 sequences, 6 sequences failed due to complete darkness, while 4 sequences encountered tracking failures caused by lighting changes and view occlusions. These limitations could potentially be addressed in the future by integrating place recognition for relocalization. Our evaluation on the Waymo Open dataset (Table III) further validates our approach, where we achieve the lowest ATE among all competing methods. This highlights the ability of our system to generalize effectively to challenging large-scale outdoor environments with complex scene geometries that typically pose substantial difficulties for monocular systems.
@@ IV-E Geometry Reconstruction Quality
. Method Metric ro-0 ro-1 ro-2 of-0 of-1 of-2 of-3 of-4 Avg. NeRF-based NICER- SLAM[58] Acc.[cm] $\downarrow$ 2.53 3.93 3.40 5.49 3.45 4.02 3.34 3.03 3.65 Comp.[cm] $\downarrow$ 3.04 4.10 3.42 6.09 4.42 4.29 4.03 3.87 4.16 Comp.Rat[%$\uparrow$ 88.75 76.61 86.10 65.19 77.84 74.51 82.01 83.98 79.37 GO-  SLAM[71] Acc.[cm] $\downarrow$ 4.60 3.31 3.97 3.05 2.74 4.61 4.32 3.91 3.81 Comp.[cm] $\downarrow$ 5.56 3.48 6.90 3.31 3.46 5.16 5.40 5.01 4.79 Comp.Rat[%$\uparrow$ 73.35 82.86 74.23 82.56 86.19 75.76 72.63 76.61 78.00 HI- SLAM[36] Acc.[cm] $\downarrow$ 3.21 3.74 3.16 3.87 2.60 4.62 4.25 3.53 3.62 Comp.[cm] $\downarrow$ 3.25 3.08 4.09 5.29 8.83 4.42 4.06 3.72 4.59 Comp.Rat[%$\uparrow$ 86.99 87.19 80.82 72.55 72.44 80.90 81.04 82.88 80.60 3DGS-based GLORIE SLAM[59] Acc.[cm] $\downarrow$ 2.84 3.07 3.05 2.98 2.06 3.32 3.34 2.92 2.96 Comp.[cm] $\downarrow$ 4.65 3.55 3.64 2.39 3.43 4.54 4.57 4.78 3.95 Comp.Rat[%$\uparrow$ 81.96 85.78 84.50 88.82 85.07 82.09 80.41 81.04 83.72 Splat- SLAM[33] Acc.[cm] $\downarrow$ 1.99 1.91 2.06 3.96 2.03 3.45 2.15 1.89 2.43 Comp.[cm] $\downarrow$ 3.78 3.38 3.34 2.75 3.33 4.36 3.96 4.25 3.64 Comp.Rat[%$\uparrow$ 85.47 86.88 86.12 87.32 85.17 81.37 82.25 82.95 84.69 Ours Acc.[cm] $\downarrow$ 1.35 1.40 1.87 1.40 1.18 1.94 1.70 1.70 1.57 Comp.[cm] $\downarrow$ 3.33 3.27 3.66 2.07 3.23 4.29 3.84 4.26 3.49 Comp.Rat[%$\uparrow$ 87.45 85.91 86.13 89.41 85.63 81.73 82.52 83.23 85.25
@@ IV-G Ablation Study
[…]
@@
TABLE VIII: Ablation on the choice of depth prior models, evaluating rendering quality, reconstruction accuracy, and inference time on the Replica dataset.
@@
TABLE IX: Ablation study on the progressive improvement in trajectory accuracy on the Replica dataset, averaged over 8 sequences. From left to right, each stage refines the pose estimation based on the previous stage.
@@
ATE $[cm]\downarrow$ 0.42 0.33 0.32 0.26
TABLE X: Ablation study on the impact of the different proposed modules on the reconstruction performance on the Replica dataset, averaged over 8 sequences.
@@
Monocular Prior Integration We first evaluate the depth estimation accuracy improved by our different modules on the Replica dataset, as depth accuracy is crucial for scene reconstruction quality. Table VII quantifies the effectiveness of different approaches. Comparing Prior (one scale) with Prior (scale grid) demonstrates that our grid-based scale alignment significantly outperforms the single-scale alignment from HI-SLAM [36], better addressing inherent scale distortions in monocular depth priors. Further analysis compares two depth estimation approaches: BA estimate (using BA alone) and BA with JDSA (using interleaved BA and JDSA optimization). The results confirm that incorporating JDSA with alternating optimization outperforms BA alone, validating our depth prior integration strategy. The final rendered depth from our Gaussian map achieves the highest accuracy, validating the effectiveness of our complete pipeline. We further investigate the incorporation of alternative depth prior predictors, including Metric3D [77] and Depth Anything (DA) V2 [78], as shown in Table VIII. The results demonstrate that our method is compatible with these alternatives. However, our chosen prior, OmniData, provides a more efficient trade-off between performance and computational cost.
@@
Trajectory Accuracy Table IX demonstrates the progressive improvement in pose estimation accuracy on the Replica dataset through our system pipeline. Starting with initial estimates from the online tracking module, accuracy is first enhanced through online PGBA based loop closing. A subsequent full BA further refines these results, with the final joint pose and 3DGS map refinement achieving the highest trajectory accuracy. This systematic improvement across stages validates the effectiveness of our hierarchical optimization approach.
Component Analysis To evaluate key design components, we conduct ablation studies by removing individual components. Table X confirms each module’s contribution to system performance. The grid-based scale alignment proves crucial, as its removal significantly degrades reconstruction accuracy. Similarly, the unbiased depth rendering enhances both rendering quality and geometric accuracy, with qualitative comparison shown in Fig. 10. While the normal loss slightly affects appearance metrics, it substantially improves geometry quality. The final joint pose and map refinement further enhances accuracy through improved global consistency.
@@ IV-H Performance Analysis
Figure 11: Map size evolution over frame index for the Scene0000 sequence in the ScanNet dataset. The Gaussian map size increases as new areas are explored but stabilizes over time due to our pruning strategy, which is triggered every 150 iterations at the peaks of the growth curve shown in the figure.
Our system achieves real-time performance with the online tracking, loop closing and mapping operating at 22 frames per second (FPS) on the Replica dataset, and takes only 12 seconds for offline refinement. On the ScanNet dataset, despite more rapid camera movements, it maintains robust performance online above 10 FPS, with offline refinement taking a few minutes due to a higher number of optimization iterations. Fig. 11 illustrates the map size evolution for sequence Scene0000. While the Gaussian count initially grows during new area exploration, our efficient map pruning strategy stabilizes the map size by eliminating redundant Gaussians in revisited regions. This demonstrates our system’s scalability and efficiency in managing large-scale environments.
Figure 12: Evaluation on self-collected data with the robot navigating through a large factory hall. Our robot (a) is equipped with stereo cameras. Our system operates solely on the monocular input from the left camera and reconstructs the scene along with the estimated camera trajectory (b). The rendered color and depth maps (c) achieve comparable visual quality while exhibiting significantly better geometry and free of floaters that are present in the baseline DROID-SLAM + 3DGS.
@@ IV-I Evaluation on Self-Collected Robot Data
We evaluate our system on a self-collected dataset captured by our robotic platform. Figure 12 shows the platform and the results in a large factory hall. The robot is equipped with stereo cameras. However, only the monocular input from the left camera is used for this experiment. The recorded sequence has intotal 4073 frames with a duration of 6 minutes and 52 seconds. We compare our rendering results with the baseline method DROID-SLAM + 3DGS, where the estimated camera poses and point cloud from DROID-SLAM serve as input and used to initialize the Gaussians. For our system, it completes online stage including tracking, loop closuing, and mapping in 5 minutes and 20 seconds, followed by 3 minutes and 32 seconds for the offline refinement stage. In contrast, the baseline requires 4 minutes and 24 seconds for DROID-SLAM and an additional 15 minutes and 2 seconds for 3DGS mapping. Overall, our method takes only about half the runtime of the baseline while achieving comparable rendering quality and significantly better geometry, without the floaters observed in the baseline results.
@@ IV-J Extention to Semantic Reconstruction
TABLE XI: Semantic reconstruction results evaluated by mIoU metric on 4 sequences of the Replica dataset.
Method ro-0 ro-1 ro-2 of-0 Avg.mIoU[%$\uparrow$
@@
Tab. XI presents quantitative results comparing our approach against both RGB-D and RGB-only semantic SLAM methods. Our system achieves competitive performance compared to state-of-the-art RGB-D methods, while significantly outperforming the RGB-only baseline HI-SLAM [36] by a margin of 14.93%. This substantial improvement can be attributed to two key factors: (1) our more accurate geometry reconstruction provides better surface boundaries for semantic label assignment, and (2) the explicit 3DGS representation allows for sharper semantic boundaries compared to implicit NeRF-based approaches that often struggle with object delineation in complex scenes or regions with small objects. Fig. 13 demonstrates qualitative results: indoor scene from the Replica dataset with ground truth semantic labels, and outdoor driving scene from the Waymo Open dataset using Mask2Former [83] predictions as semantic inputs.

### v3

@@ Abstract
We present HI-SLAM2, a geometry-aware Gaussian SLAM system that achieves fast and accurate monocular scene reconstruction using only RGB input. Existing Neural SLAM or 3DGS-based SLAM methods often trade off between rendering quality and geometry accuracy, our research demonstrates that both can be achieved simultaneously with RGB input alone. The key idea of our approach is to enhance the ability for geometry estimation by combining easy-to-obtain monocular priors with learning-based dense SLAM, and then using 3D Gaussian splatting as our core map representation to efficiently model the scene. Upon loop closure, our method ensures on-the-fly global consistency through efficient pose graph bundle adjustment and instant map updates by explicitly deforming the 3D Gaussian units based on anchored keyframe updates. Furthermore, we introduce a grid-based scale alignment strategy to maintain improved scale consistency in prior depths for finer depth details. Through extensive experiments on Replica, ScanNet, Waymo Open, ETH3D SLAM and ScanNet++ datasets, we demonstrate significant improvements over existing Neural SLAM methods and even surpass RGB-D-based methods in both reconstruction and rendering quality. The project page and source code are available at https://hi-slam2.github.io/.
@@ III-A Comparison to HI-SLAM
  • Hierarchical Optimization: Our system employs a multi-stage optimization pipeline that includes online tracking with local BA, online loop closure with $Sim(3)$-based PGBA, and global full BA. Last but not least, the joint refinement of both Gaussian map parameters and camera poses, which couples the tracking frontend with the mapping backend more tightly. In contrast, HI-SLAM only performs pose optimization in the tracking frontend, which can lead to inconsistencies between estimated poses and map geometry. This hierarchical approach reduces the Absolute Trajectory Error (ATE) by 29.3% compared to online tracking alone (Table X) on the Replica dataset, yielding a globally more consistent reconstruction.
@@ IV-D Camera Tracking Accuracy
We evaluate the camera tracking accuracy of our system against state-of-the-art dense visual SLAM methods on indoor Replica, ScanNet, ETH3D SLAM datasets and outdoor Waymo datasets, including comparisons with RGB-D based approaches. The ATE results in Tables I and II demonstrate the superior tracking accuracy of our system. RGB-D methods SplatTAM [28] and MonoGS [38], despite having access to depth measurements for map-based tracking, achieve lower accuracy than hybrid approaches. Our system, along with Splat-SLAM [53], represents the class of hybrid methods that effectively combine dense SLAM with deep learning foundations. The global BA of DROID-SLAM [63] was enabled in all experiments. While it employs global BA, our additional global pose and map joint refinement further improves tracking accuracy beyond the baseline method. We further compare our system on the ETH3D SLAM dataset with state-of-the-art sparse and dense methods, including ORB-SLAM3 [7], DPVO [64], DPV-SLAM [36], and MASt3R-SLAM [42]. Figure 6 illustrates the cumulative success curves based on ATE thresholds and ATE statistics. As none of the methods can successfully track all sequences, we use the Area Under the Curve (AUC) metric to assess both accuracy and robustness with an upper ATE threshold of 0.5m. Our system achieves the highest AUC among all methods and lowest mean and median ATE. Out of total 61 sequences, 6 sequences failed due to complete darkness, while 4 sequences encountered tracking failures caused by lighting changes and view occlusions. These limitations could potentially be addressed in the future by integrating place recognition for relocalization to reduce the standard deviation of ATE. Our evaluation on the Waymo Open dataset (Table III) further validates our approach, where we achieve the lowest ATE among all competing methods. This highlights the ability of our system to generalize effectively to challenging large-scale outdoor environments with complex scene geometries that typically pose substantial difficulties for monocular systems.
@@ IV-E Geometry Reconstruction Quality
. Method Metric ro-0 ro-1 ro-2 of-0 of-1 of-2 of-3 of-4 Avg. NeRF-based NICER- SLAM[85] Acc.[cm] $\downarrow$ 2.53 3.93 3.40 5.49 3.45 4.02 3.34 3.03 3.65 Comp.[cm] $\downarrow$ 3.04 4.10 3.42 6.09 4.42 4.29 4.03 3.87 4.16 Comp.Rat[$\uparrow$ 88.75 76.61 86.10 65.19 77.84 74.51 82.01 83.98 79.37 GO-  SLAM[81] Acc.[cm] $\downarrow$ 4.60 3.31 3.97 3.05 2.74 4.61 4.32 3.91 3.81 Comp.[cm] $\downarrow$ 5.56 3.48 6.90 3.31 3.46 5.16 5.40 5.01 4.79 Comp.Rat[$\uparrow$ 73.35 82.86 74.23 82.56 86.19 75.76 72.63 76.61 78.00 HI- SLAM[79] Acc.[cm] $\downarrow$ 3.21 3.74 3.16 3.87 2.60 4.62 4.25 3.53 3.62 Comp.[cm] $\downarrow$ 3.25 3.08 4.09 5.29 8.83 4.42 4.06 3.72 4.59 Comp.Rat[$\uparrow$ 86.99 87.19 80.82 72.55 72.44 80.90 81.04 82.88 80.60 3DGS-based GLORIE SLAM[77] Acc.[cm] $\downarrow$ 2.84 3.07 3.05 2.98 2.06 3.32 3.34 2.92 2.96 Comp.[cm] $\downarrow$ 4.65 3.55 3.64 2.39 3.43 4.54 4.57 4.78 3.95 Comp.Rat[$\uparrow$ 81.96 85.78 84.50 88.82 85.07 82.09 80.41 81.04 83.72 Splat- SLAM[53] Acc.[cm] $\downarrow$ 1.99 1.91 2.06 3.96 2.03 3.45 2.15 1.89 2.43 Comp.[cm] $\downarrow$ 3.78 3.38 3.34 2.75 3.33 4.36 3.96 4.25 3.64 Comp.Rat[$\uparrow$ 85.47 86.88 86.12 87.32 85.17 81.37 82.25 82.95 84.69 Ours Acc.[cm] $\downarrow$ 1.35 1.40 1.87 1.40 1.18 1.94 1.70 1.70 1.57 Comp.[cm] $\downarrow$ 3.33 3.27 3.66 2.07 3.23 4.29 3.84 4.26 3.49 Comp.Rat[$\uparrow$ 87.45 85.91 86.13 89.41 85.63 81.73 82.52 83.23 85.25
@@ IV-G Ablation Study
Abs Diff $m]\downarrow$ Abs Rel $\%$$\downarrow$ Sq Rel $\%$$\downarrow$ RMSE $m]\downarrow$ $\delta<$1.05 $\%$$\uparrow$ $\delta<$1.25 $\%$$\uparrow$
@@
TABLE VIII: Ablation study on the choice of depth prior models, evaluating rendering quality, reconstruction accuracy, and inference time per frame on the Replica dataset.
@@
TABLE IX: Ablation study on the choice of normal prior models.
Ours + EESNU 38.60 1.70 3.64 84.61 5
Ours + DSINE 38.52 1.62 3.55 85.02 37
Ours + OmniData 38.71 1.57 3.49 85.25 6
TABLE X: Ablation study on the progressive improvement in trajectory accuracy on the Replica dataset, averaged over 8 sequences. From left to right, each stage refines the pose estimation based on the previous stage.
@@
ATE $cm]\downarrow$ 0.42 0.33 0.32 0.26
TABLE XI: Ablation study on the impact of the different proposed modules on the reconstruction performance on the Replica dataset, averaged over 8 sequences.
@@
Monocular Prior Integration We first evaluate the depth estimation accuracy improved by our different modules on the Replica dataset, as depth accuracy is crucial for scene reconstruction quality. Table VII quantifies the effectiveness of different approaches. Comparing Prior (one scale) with Prior (scale grid) demonstrates that our grid-based scale alignment significantly outperforms the single-scale alignment from HI-SLAM [79], better addressing inherent scale distortions in monocular depth priors. Further analysis compares two depth estimation approaches: BA estimate (using BA alone) and BA with JDSA (using interleaved BA and JDSA optimization). The results confirm that incorporating JDSA with alternating optimization outperforms BA alone, validating our depth prior integration strategy. The final rendered depth from our Gaussian map achieves the highest accuracy, validating the effectiveness of our complete pipeline.
We further investigate the incorporation of alternative depth and normal prior predictors. For depth priors, including Metric3D [73], ZoeDepth [5], and Depth Anything (DA) V2 [70], we present results in Table VIII. Notably, ZoeDepth and DA V2 were not trained on any Replica images, yet still achieve comparable performance, indicating that our method does not rely on indirect biases from the training data. For normal priors, we evaluate EESNU [2] and DSINE [3], as shown in Table IX. The results demonstrate that our method remains compatible with these alternatives. Nevertheless, our chosen prior, OmniData, offers a more efficient trade-off between performance and computational cost. This choice also aligns with the baseline methods, which likewise employ OmniData as the geometry prior.
@@
Trajectory Accuracy Table X demonstrates the progressive improvement in pose estimation accuracy on the Replica dataset through our system pipeline. Starting with initial estimates from the online tracking module, accuracy is first enhanced through online PGBA based loop closing. A subsequent full BA further refines these results, with the final joint pose and 3DGS map refinement achieving the highest trajectory accuracy. This systematic improvement across stages validates the effectiveness of our hierarchical optimization approach.
Component Analysis To evaluate key design components, we conduct ablation studies by removing individual components. Table XI confirms each module’s contribution to system performance. The grid-based scale alignment proves crucial, as its removal significantly degrades reconstruction accuracy. Similarly, the unbiased depth rendering enhances both rendering quality and geometric accuracy, with qualitative comparison shown in Fig. 10. While the normal loss slightly affects appearance metrics, it substantially improves geometry quality. The final joint pose and map refinement further enhances accuracy through improved global consistency.
@@ IV-H Performance Analysis
Figure 11: Evolution of map size, GPU memory usage, and system speed over frame index for the Scene0000 sequence from ScanNet, evaluated on an RTX 4090 GPU. The number of Gaussians grows as new areas are explored and stabilizes with the pruning strategy. System speed denotes the processing frame rate for online tracking, loop closing, and mapping. Allocated and reserved GPU memory usage are both reported to provide resource analysis.
Our system achieves real-time performance with the online tracking, loop closing and mapping operating at 22 frames per second (FPS) on the Replica dataset, and takes only 12 seconds for offline refinement. On the ScanNet dataset, despite more rapid camera movements, it maintains robust performance online above 10 FPS, with offline refinement taking a few minutes due to a higher number of optimization iterations. Fig. 11 illustrates the map size evolution, GPU memory consumption, and processing speed for sequence Scene0000. While the Gaussian count initially grows during new area exploration, our efficient map pruning strategy stabilizes the map size by eliminating redundant Gaussians in revisited regions. The memory analysis reveals that both allocated and reserved GPU memory usage remain stable throughout the sequence, with periodic fluctuations corresponding to pruning cycles. Reserved memory gradually increases as additional buffers are allocated to store keyframe states. The processing frame rate decreases during rapid motion due to more frequent keyframe insertion, and increases during stable motion, remaining consistent across different phases. This demonstrates our system’s scalability and efficiency in managing large-scale environments while maintaining predictable resource consumption.
Figure 12: Evaluation on self-collected data with the robot navigating through a large factory hall. Our robot (a) is equipped with stereo cameras. Our system operates solely on the monocular input from the left camera and reconstructs the scene along with the estimated camera trajectory (b). The rendered color and depth maps (c) achieve comparable visual quality while exhibiting significantly better geometry and free of floaters. In contrast, the baseline DROID-SLAM + 3DGS (d) suffers from geometric artifacts and floating points. We refer readers to our supplementary video, which showcases novel view renderings on out-of-sequence viewpoints.
@@ IV-I Evaluation on Self-Collected Robot Data
Figure 13: Qualitative comparison of the estimated robot trajectory in the factory hall.
We evaluate our system on a self-collected dataset captured by our robotic platform. Figure 12 shows the platform and the results in a large factory hall. The robot is equipped with stereo cameras. However, only the monocular input from the left camera is used for this experiment. The recorded sequence has intotal 4073 frames with a duration of 6 minutes and 52 seconds. We compare our rendering results with the baseline method DROID-SLAM + 3DGS, where the estimated camera poses and point cloud from DROID-SLAM serve as input and used to initialize the Gaussians. For our system, it completes online stage including tracking, loop closuing, and mapping in 5 minutes and 20 seconds, followed by 3 minutes and 32 seconds for the offline refinement stage. In contrast, the baseline requires 4 minutes and 24 seconds for DROID-SLAM and an additional 15 minutes and 2 seconds for 3DGS mapping. Overall, our method takes only about half the runtime of the baseline while achieving comparable rendering quality and significantly better geometry, without the floaters observed in the baseline results. Furthermore, we evaluate the trajectory accuracy using the centimeter-accurate photogrammetric reference described in [50], as shown in Fig. 13. While MASt3R-SLAM [42] struggles to track camera poses due to poor generalizability to the new environment, our system maintains stable tracking and successfully detects loop closures with the highest accuracy.
@@ IV-J Extention to Semantic Reconstruction
TABLE XII: Semantic reconstruction results evaluated by mIoU metric on 4 sequences of the Replica dataset.
Method ro-0 ro-1 ro-2 of-0 Avg.mIoU[$\uparrow$
@@
Tab. XII presents quantitative results comparing our approach against recent RGB-D and RGB-only semantic SLAM methods, including concurrent work Hier-SLAM [33]. Our system achieves competitive performance compared to state-of-the-art RGB-D methods, while significantly outperforming the RGB-only baseline HI-SLAM [79] by a margin of 14.93%. This substantial improvement can be attributed to two key factors: (1) our more accurate geometry reconstruction provides better surface boundaries for semantic label assignment, and (2) the explicit 3DGS representation allows for sharper semantic boundaries compared to implicit NeRF-based approaches that often struggle with object delineation in complex scenes or regions with small objects. Fig. 14 demonstrates qualitative results: indoor scene from the Replica dataset with ground truth semantic labels, and outdoor driving scene from the Waymo Open dataset using Mask2Former [9] predictions as semantic inputs.